# Exotic Particles at the DUNE Near Detector from Charged Pion Scattering

### v1

@@ I Introduction
In the search for light, weakly-coupled extensions to the Standard Model, fixed-target experiments have proved to be an extremely useful approach (see Ref. [1] and references therein). Trading energy for luminosity compared to colliding-beam experiments, these $\mathcal{O}(10-100)\ {\rm GeV}$ proton and electron beams can deliver Avogadro’s number of beam particles on target, allowing us to search for a wide variety of low-mass particles beyond the Standard Model (BSM) with small production cross sections. In particular, proton beams produce copious amounts of secondary particles, including neutral mesons and muons, which can generate BSM particles through rare decays, mixing, or bremsstrahlung. Recent studies [2, 3, 4, 5] have demonstrated that as part of the upcoming DUNE program [6, 7, 8, 9], the neutral meson channel yields excellent sensitivity to millicharged particles (MCPs) (denoted $\chi$) or heavy axion-like particles (ALPs) coupled to QCD (denoted $a$).
@@
However, previous analyses have overlooked an additional production mechanism for these new particles, namely charged pion scattering. At DUNE, the 120 GeV proton beam creates $\sim 6.5\,\pi^{\pm}$ for every proton on target (POT) compared to $\sim 3.5\,\pi^{0}$ per POT [10]. As a charged pion traverses the target, it can scatter off the target nuclei and produce BSM particles through bremsstrahlung-like processes, which may be described in the framework of chiral perturbation theory as long as the BSM particles have masses well below $4\pi f_{\pi}\sim 1\ {\rm GeV}$, where $f_{\pi}=93\ {\rm MeV}$ is the pion decay constant. This production channel was first considered for dark photon production at DarkQuest [11], where it was shown that charged pion bremsstrahlung generated comparable event rates to the Drell-Yan production channel for GeV-scale dark photons.
@@
In this paper, we follow the logic of Ref. [11] to search for MCPs and heavy ALPs at DUNE through secondary charged pion scattering. The MCP production channel is very similar to the previously-studied dark photon production, involving an off-shell virtual photon generating a $\chi\bar{\chi}$ pair instead of an on-shell dark photon decaying to visible states, and we will show that charged pion scattering does indeed yield modest gains in sensitivity for GeV-scale $\chi$. For heavy axions, QCD-coupled axion production from charged pions dominates the meson mixing production channel except in narrow resonance windows around the neutral meson masses, and is comparable to the expected rate from kaon decays [5]. At high axion masses, charged pion scattering provides additional sensitivity at large couplings all the way up to the regime of validity of chiral perturbation theory.
@@
This paper is organized as follows. In Sec. II, we briefly describe the experimental design of the DUNE Near Detector (ND) Complex located at Fermilab, including additional detector components proposed for detecting MCPs. In Sec. III, we give an overview of our simulation pipeline that we use to model the production and detection of BSM particles from charged pion scattering. In Sec. IV, we describe our MCP model and calculate the additional sensitivity gained at DUNE ND at high MCP masses from charged-pion scattering, compared to the production modes previously considered in Refs. [2, 3]. In Sec. V, we do the same for the heavy axion model of Ref. [4]. We conclude in Sec. VI.
@@ II DUNE as a BSM detector
Figure 1: A simple schematic of the experimental setup for MCP (left) and ALP (right) searches. Note that the DUNE ND Complex contains ArgonCube and the Multi-Purpose Detector, which are the detectors used for our ALP search. Additionally, for the sake of our MCP search, one could imagine installing scintillator arrays and PMTs inside the Dune ND Complex inspired by the FerMINI proposal [2], described in more detail in Sec. IV.2.
DUNE is primarily designed to produce neutrinos, but its 120 GeV proton beam can also be used as a source for BSM particles, which can be detected at the DUNE Near Detector (ND) Complex. The LBNF-DUNE beam is a 120 GeV proton beam incident on a graphite target, with an expected luminosity of $1.47\times 10^{21}$ protons on target (POT) each year, and a total of $1.47\times 10^{22}$ POT over a 10-year run [12, 6, 7, 8, 9]. The ND Complex is located 574 meters down the beam line from the target, and will contain two argon detectors: ArgonCube, a 50 ton liquid argon time projection chamber (LArTPC), and the Multi-Purpose Detector (MPD), a gaseous argon time projection chamber designed to detect muons and other particles that are not stopped in ArgonCube. ArgonCube is 7 m in width, 3 m in height, and 5 m in length along the beamline. The MPD is situated directly downstream from ArgonCube and has a cylindrical volume that is 5 m in diameter and 5 m in height. Both ArgonCube and the MPD are sensitive to diphoton and hadron signatures by reconstructing the invariant mass and direction of the event, and these will be the primary decay modes of the heavy axions we consider.
To search for MCPs, one may also imagine adding scintillator arrays in the Near Detector Complex, as was proposed in [2]. To make a direct comparison with Ref. [2], we consider the same setup: three stacks of scintillator arrays in a 1 m $\times$ 1 m detection area, coupled to photomultiplier tubes (PMTs) to detect charged particles. The distinctive signature of a MCP would be soft ionizations with yield below that expected from a minimum-ionizing particle of charge $|Q|=1$.
@@ III.1 Charged Pion Production
When the 120 GeV DUNE proton beam collides with the target, it produces a multitude of SM mesons, including $\pi,\eta,\eta^{\prime}$ as well as heavier mesons. While mesons are produced during several nuclear interaction lengths, in order to provide a conservative estimate of the reach and to make a direct comparison with previous studies [10, 2, 3, 4, 11], we only consider the mesons produced within the first thin layer of the target.
@@
Figure 2: The energy distribution of charged pions from a Pythia 8.2 simulation of a 120 GeV proton hitting a proton target. The pions were created within the first thin slice of the target. We rescale this distribution to match the nucleon content of the DUNE graphite target. This distribution was adapted from the analysis of [10].
@@ III.2 BSM Particle Flux
Exotic particles may be created from pion-nucleon scattering, $\pi^{\pm}N\to\pi^{\pm}N^{\prime}X$, where we will later consider $X=a$ or $\chi\bar{\chi}$. [^2] Axions can also be produced through pion conversion, $\pi N\to aN$, though we will discuss in Sec. V that this process is (surprisingly) subdominant to the 3-body process. For sufficiently small momentum transfers, we can describe the scattering process using chiral perturbation theory. In this approximation, the pions would scatter incoherently off the nucleons contained within the target’s nuclei. Typically, staying within the regime of validity of chiral perturbation theory limits the mass of the BSM particles to $\lesssim 1\ {\rm GeV}$, though for MCPs in particular, the production rate remains significant up to about 5 GeV. A full treatment of this process would require matching onto the perturbative QCD regime, which is beyond the scope of this work but has significant synergy with the DUNE neutrino cross section program [13].
@@
where $E_{\pi}$ is the energy of the incoming pion, $M_{n}$ is the nucleon mass, $d\Pi$ is the 3- or 4-body phase space measure, $Z=6$ is the atomic number of the graphite target, $A=12$ is the mass number, and $\langle|M^{(p,n)}|^{2}\rangle$ are the spin-averaged matrix elements for $\pi^{\pm}$ scattering off protons or neutrons, respectively. [^3] There is a subtlety in treating the target mass in the cross section calculation. Here, we have treated the pions as scattering off individual nucleons, but the full center-of-mass energy (and hence the kinematic threshold) of the scattering process involves the nucleus mass. To treat the target mass properly, we should apply a target mass correction [14], but in this work we have chosen to take the conservative approach (with the lowest possible kinematic threshold) and use the nucleon mass only. For both axion and MCP production, we perform a Monte Carlo integral of the cross section using MadGraph5 [15] to generate events and implement a momentum cutoff at $|t|=-(p^{\prime}_{n}-p_{n})^{2}\leq(4\pi f_{\pi})^{2}$ as a phase space cut, where $p_{n}$ and $p^{\prime}_{n}$ are the initial and final nucleon 4-vectors, respectively. Because MCP production proceeds through a virtual photon, low-mass MCPs will be dominantly produced in the region of small photon virtuality, where the soft singularity of QED makes the Monte Carlo sampling unreliable. Consequently, we will focus only on MCPs above about 400 MeV, and we will use the closely-related study of dark photon production in [11] to argue that there are limited sensitivity gains for lighter MCPs.
@@ IV Millicharged Particles
To make contact with previous studies of MCPs [2, 3], we consider a model where a fermionic MCP $\chi$ with mass $m_{\chi}$ has electric charge $Q=e\epsilon$, with $e=\sqrt{4\pi\alpha}$ the QED gauge coupling and $\epsilon\ll 1$. The MCP therefore couples to the SM photon $A_{\mu}$ through the Lagrangian
@@
where $\mathcal{L}_{\pi N}$ and $\mathcal{L}_{\pi NA}$ are pion-nucleon and pion-nucleon-photon Lagrangians described by chiral perturbation theory. Following Refs. [16, 17, 18, 19, 20, 11], pion-nucleon interactions can be described by
@@
where $p$ and $n$ are the proton and neutron spinors, $\pi^{0}$ and $\pi^{\pm}$ are the pion fields, $f_{\pi}=93$ MeV is the pion decay constant, and $g_{A}=1.27$ is the axial coupling determined from neutron beta decay [21]. Similarly, the pion-nucleon-photon interactions are described by
@@
Our signal process is the scattering process $\pi^{\pm}N\to\pi^{\pm}N\chi\bar{\chi}$, where the $\chi\bar{\chi}$ pair is produced from a virtual photon radiated by the charged pion or the proton. The leading-order Feynman diagrams for MCP production from $\pi^{-}p$ scattering are shown in Fig. 3. Note that in order to preserve the Ward identity, we must sum over all diagrams and cannot claim that a subset are parametrically suppressed compared to the rest; the usual intuition that bremsstrahlung is suppressed for radiation by heavy particles does not hold in the non-renormalizable chiral Lagrangian [11]. A similar set of diagrams related by crossing and isospin symmetry exist for $\pi^{-}n$, $\pi^{+}p$, and $\pi^{+}n$ scattering. To obtain the full cross section, we must sum incoherently over all four channels.
@@ IV.1 MCP Flux
Given the large multiplicity of Feynman diagrams contributing to $\pi N\to\pi N\chi\bar{\chi}$ and the 4-body final state, we use MadGraph to calculate the cross section with a momentum cutoff $|t|<(4\pi f_{\pi})^{2}$ that ensures our cross sections are reliable within the regime of validity of chiral perturbation theory. As shown in Fig. 4, the cross section is only weakly dependent on incident pion energy above the kinematic threshold for $\chi\bar{\chi}$ production.
Figure 4: The total inclusive cross section for the scattering process $\pi N\to\pi N\chi\bar{\chi}$, summed over all charged pion and nucleon channels as in Eq. (2), as a function of incoming pion energy $E_{\pi}$ for various MCP masses $m_{\chi}$.
@@
To model how many MCPs can reach the detector, we implemented an angular cut on MCP trajectories as depicted in Fig. 1. For an expected detector area of $1\ {\rm m}^{2}$ at a distance of 574 m from the target, the angular acceptance is $\theta_{\rm acc.}=8.71\times 10^{-4}\;\text{rad}$. The expected flux of MCPs as a function of MCP mass is shown in Fig. 5. For $m_{\chi}>400\ {\rm MeV}$, we see that the MCP flux from charged pion production exceeds that of Drell-Yan production all the way up to the kinematic threshold of the 120 GeV beam. Interestingly, this is true even for the conservative assumption of the target mass equal to the bare nucleon mass, while for dark photon production in Ref. [11], the relative benefit of charged pion scattering is much smaller. As alluded to earlier, MadGraph becomes unreliable for small MCP masses as the virtuality of the photon becomes small, so we do not show the flux for $m_{\chi}<400\ {\rm MeV}$. That said, based on the results of Ref. [11] for similar kinematics, we strongly expect that the flux will be subdominant to meson decays in this mass range and will not yield any additional gains in sensitivity.
Figure 5: Number of MCP $N_{\chi}$ incident at the DUNE ND, for $10^{21}$ POT and $\epsilon=10^{-3}$. Our production from meson decays are consistent with Ref. [2], but charged pion scattering exceeds Drell-Yan scattering up to highest masses ($\sim 5\ {\rm GeV}$) allowed by the chiral cutoff, and is thus the dominant production mode at high masses.

### v2

@@ I Introduction
In the search for light, weakly-coupled extensions to the Standard Model, fixed-target experiments have proved to be an extremely useful approach (see Ref. Gori et al. (2022) and references therein). Trading energy for luminosity compared to colliding-beam experiments, these $\mathcal{O}(10-100)\ {\rm GeV}$ proton and electron beams can deliver Avogadro’s number of beam particles on target, allowing us to search for a wide variety of low-mass particles beyond the Standard Model (BSM) with small production cross sections. In particular, proton beams produce copious amounts of secondary particles, including neutral mesons and muons, which can generate BSM particles through rare decays, mixing, or bremsstrahlung. Recent studies Kelly and Tsai (2019); Harnik et al. (2019); Kelly et al. (2021); Berger and Putnam (2024) have demonstrated that as part of the upcoming DUNE program Acciarri et al. (2015); Strait et al. (2016); Acciarri et al. (2016a, b), the neutral meson channel yields excellent sensitivity to millicharged particles (MCPs) (denoted $\chi$) or heavy axion-like particles (ALPs) coupled to QCD (denoted $a$).
@@
However, previous analyses have overlooked an additional production mechanism for these new particles, namely charged pion scattering. At DUNE, the 120 GeV proton beam creates $\sim 6.5\,\pi^{\pm}$ for every proton on target (POT) compared to $\sim 3.5\,\pi^{0}$ per POT Berlin et al. (2018). As a charged pion traverses the target, it can scatter off the target nuclei and produce BSM particles through bremsstrahlung-like processes, which may be described in the framework of chiral perturbation theory as long as the BSM particles have masses well below $4\pi f_{\pi}\sim 1.2\ {\rm GeV}$, where $f_{\pi}=93\ {\rm MeV}$ is the pion decay constant. This production channel was first considered for dark photon production at DarkQuest Curtin et al. (2023), where it was shown that charged pion bremsstrahlung generated comparable event rates to the Drell-Yan production channel for GeV-scale dark photons.
@@
In this paper, we follow the logic of Ref. Curtin et al. (2023) to search for MCPs and heavy ALPs at DUNE through secondary charged pion scattering. The MCP production channel is very similar to the previously-studied dark photon production, involving an off-shell virtual photon generating a $\chi\bar{\chi}$ pair instead of an on-shell dark photon decaying to visible states, and we will show that charged pion scattering does indeed yield modest gains in sensitivity for GeV-scale $\chi$. For heavy axions, QCD-coupled axion production from charged pions dominates the meson mixing production channel except in narrow resonance windows around the neutral meson masses, and is comparable to the expected rate from kaon decays Berger and Putnam (2024). At high axion masses, charged pion scattering provides additional sensitivity at large couplings all the way up to the regime of validity of chiral perturbation theory.
@@
This paper is organized as follows. In Sec. II, we briefly describe the experimental design of the DUNE Near Detector (ND) Complex located at Fermilab, including additional detector components proposed for detecting MCPs. In Sec. III, we give an overview of our simulation pipeline that we use to model the production and detection of BSM particles from charged pion scattering. In Sec. IV, we describe our MCP model and calculate the additional sensitivity gained at DUNE ND at high MCP masses from charged-pion scattering, compared to the production modes previously considered in Refs. Kelly and Tsai (2019); Harnik et al. (2019). In Sec. V, we do the same for the heavy axion model of Ref. Kelly et al. (2021). We conclude in Sec. VI.
@@ II DUNE as a BSM detector
Figure 1: A simple schematic of the experimental setup for MCP (left) and ALP (right) searches. Note that the DUNE ND Complex contains ArgonCube and the Multi-Purpose Detector, which are the detectors used for our ALP search. Additionally, for the sake of our MCP search, one could imagine installing scintillator arrays and PMTs inside the Dune ND Complex inspired by the FerMINI proposal Kelly and Tsai (2019), described in more detail in Sec. IV.2.
DUNE is primarily designed to produce neutrinos, but its 120 GeV proton beam can also be used as a source for BSM particles, which can be detected at the DUNE Near Detector (ND) Complex. The LBNF-DUNE beam is a 120 GeV proton beam incident on a graphite target, with an expected luminosity of $1.47\times 10^{21}$ protons on target (POT) each year, and a total of $1.47\times 10^{22}$ POT over a 10-year run Berryman et al. (2020); Acciarri et al. (2015); Strait et al. (2016); Acciarri et al. (2016a, b). The ND Complex is located 574 meters down the beam line from the target, and will contain two argon detectors: ArgonCube, a 50 ton liquid argon time projection chamber (LArTPC), and the Multi-Purpose Detector (MPD), a gaseous argon time projection chamber designed to detect muons and other particles that are not stopped in ArgonCube. ArgonCube is 7 m in width, 3 m in height, and 5 m in length along the beamline. The MPD is situated directly downstream from ArgonCube and has a cylindrical volume that is 5 m in diameter and 5 m in height. Both ArgonCube and the MPD are sensitive to diphoton and hadron signatures by reconstructing the invariant mass and direction of the event, and these will be the primary decay modes of the heavy axions we consider.
To search for MCPs, one may also imagine adding scintillator arrays in the Near Detector Complex, as was proposed in Kelly and Tsai (2019). To make a direct comparison with Ref. Kelly and Tsai (2019), we consider the same setup: three stacks of scintillator arrays in a 1 m $\times$ 1 m detection area, coupled to photomultiplier tubes (PMTs) to detect charged particles. The distinctive signature of a MCP would be soft ionizations with yield below that expected from a minimum-ionizing particle of charge $|Q|=1$.
@@ III.1 Charged Pion Production
When the 120 GeV DUNE proton beam collides with the target, it produces a multitude of SM mesons, including $\pi,\eta,\eta^{\prime}$ as well as heavier mesons. While mesons are produced during several nuclear interaction lengths, in order to provide a conservative estimate of the reach and to make a direct comparison with previous studies Berlin et al. (2018); Kelly and Tsai (2019); Harnik et al. (2019); Kelly et al. (2021); Curtin et al. (2023), we only consider the mesons produced within the first thin layer of the target.
@@
Figure 2: The energy distribution of charged pions from a Pythia 8.2 simulation of a 120 GeV proton hitting a proton target. The pions were created within the first thin slice of the target. We rescale this distribution to match the nucleon content of the DUNE graphite target. This distribution was adapted from the analysis of Berlin et al. (2018).
@@ III.2 BSM Particle Flux
Exotic particles may be created from pion-nucleon scattering, $\pi^{\pm}N\to\pi^{\pm}N^{\prime}X$, where we will later consider $X=a$ or $\chi\bar{\chi}$. [^2] Axions can also be produced through pion conversion, $\pi N\to aN$, though we will discuss in Sec. V that this process is (surprisingly) subdominant to the 3-body process. For sufficiently small momentum transfers, we can describe the scattering process using chiral perturbation theory. In this approximation, the pions would scatter incoherently off the nucleons contained within the target’s nuclei. Typically, staying within the regime of validity of chiral perturbation theory limits the mass of the BSM particles to $\lesssim 1\ {\rm GeV}$, though for MCPs in particular, the production rate remains significant up to about 5 GeV. A full treatment of this process would require matching onto the perturbative QCD regime, which is beyond the scope of this work but has significant synergy with the DUNE neutrino cross section program Ruso et al. (2022).
@@
where $E_{\pi}$ is the energy of the incoming pion, $M_{n}$ is the nucleon mass, $d\Pi$ is the 3- or 4-body phase space measure, $Z=6$ is the atomic number of the graphite target, $A=12$ is the mass number, and $\langle|M^{(p,n)}|^{2}\rangle$ are the spin-averaged matrix elements for $\pi^{\pm}$ scattering off protons or neutrons, respectively. [^3] There is a subtlety in treating the target mass in the cross section calculation. Here, we have treated the pions as scattering off individual nucleons, but the full center-of-mass energy (and hence the kinematic threshold) of the scattering process involves the nucleus mass. To treat the target mass properly, we should apply a target mass correction Ruiz et al. (2024), but in this work we have chosen to take the conservative approach (with the lowest possible kinematic threshold) and use the nucleon mass only. For both axion and MCP production, we perform a Monte Carlo integral of the cross section using MadGraph5 Alwall et al. (2014) to generate events. To remain within the regime of validity of leading-order chiral perturbation theory, we implement momentum cutoffs,
$t_{n}$ $\equiv-(p^{\prime}_{n}-p_{n})^{2}\leq(4\pi f_{\pi})^{2}$ (3)
$t_{\pi}$ $\equiv-(p^{\prime}_{\pi}-p_{\pi})^{2}\leq(4\pi f_{\pi})^{2}$ (4)
as phase space cuts, where $p_{n}\ (p_{\pi})$ and $p^{\prime}_{n}\ (p^{\prime}_{\pi})$ are the initial and final nucleon (pion) 4-vectors, respectively. Because MCP production proceeds through a virtual photon, low-mass MCPs will be dominantly produced in the region of small photon virtuality, where the soft singularity of QED makes the Monte Carlo sampling unreliable. Consequently, we will focus only on MCPs above about 400 MeV, and we will use the closely-related study of dark photon production in Curtin et al. (2023) to argue that there are limited sensitivity gains for lighter MCPs.
@@ IV Millicharged Particles
To make contact with previous studies of MCPs Kelly and Tsai (2019); Harnik et al. (2019), we consider a model where a fermionic MCP $\chi$ with mass $m_{\chi}$ has electric charge $Q=e\epsilon$, with $e=\sqrt{4\pi\alpha}$ the QED gauge coupling and $\epsilon\ll 1$. The MCP therefore couples to the SM photon $A_{\mu}$ through the Lagrangian
@@
where $\mathcal{L}_{\pi N}$ and $\mathcal{L}_{\pi NA}$ are pion-nucleon and pion-nucleon-photon Lagrangians described by chiral perturbation theory. Following Refs. Scherer (2003); Shin and Yun (2022); Ellis and Tang (1998); Bernard et al. (1995, 1997); Curtin et al. (2023), pion-nucleon interactions can be described by
@@
where $p$ and $n$ are the proton and neutron spinors, $\pi^{0}$ and $\pi^{\pm}$ are the pion fields, $f_{\pi}=93$ MeV is the pion decay constant, and $g_{A}=1.27$ is the axial coupling determined from neutron beta decay Czarnecki et al. (2018). Similarly, the pion-nucleon-photon interactions are described by
@@
Our signal process is the scattering process $\pi^{\pm}N\to\pi^{\pm}N\chi\bar{\chi}$, where the $\chi\bar{\chi}$ pair is produced from a virtual photon radiated by the charged pion or the proton. The leading-order Feynman diagrams for MCP production from $\pi^{-}p$ scattering are shown in Fig. 3. Note that in order to preserve the Ward identity, we must sum over all diagrams and cannot claim that a subset are parametrically suppressed compared to the rest; the usual intuition that bremsstrahlung is suppressed for radiation by heavy particles does not hold in the non-renormalizable chiral Lagrangian Curtin et al. (2023). A similar set of diagrams related by crossing and isospin symmetry exist for $\pi^{-}n$, $\pi^{+}p$, and $\pi^{+}n$ scattering. To obtain the full cross section, we must sum incoherently over all four channels.
@@ IV.1 MCP Flux
Given the large multiplicity of Feynman diagrams contributing to $\pi N\to\pi N\chi\bar{\chi}$ and the 4-body final state, we use MadGraph to calculate the cross section with momentum cutoffs given by Eqs. (3)–(4) that ensures our cross sections are reliable within the regime of validity of chiral perturbation theory. As shown in Fig. 4, the cross section is only weakly dependent on incident pion energy above the kinematic threshold for $\chi\bar{\chi}$ production.
Figure 4: The total inclusive cross section for the scattering process $\pi N\to\pi N\chi\bar{\chi}$, summed over all charged pion and nucleon channels as in Eq. (2), as a function of incoming pion energy $E_{\pi}$ for various MCP masses $m_{\chi}$. The apparent increase in the cross section at low $E_{\pi}$ for $m_{\chi}=400\ {\rm MeV}$ is an artifact of the soft singularity of QED in MadGraph, which is why we restrict our analysis to $m_{\chi}>400\ {\rm MeV}$.
@@
To model how many MCPs can reach the detector, we implemented an angular cut on MCP trajectories as depicted in Fig. 1. For an expected detector area of $1\ {\rm m}^{2}$ at a distance of 574 m from the target, the angular acceptance is $\theta_{\rm acc.}=8.71\times 10^{-4}\;\text{rad}$. The expected flux of MCPs as a function of MCP mass is shown in Fig. 5. For $m_{\chi}\gtrsim 1\ {\rm GeV}$, we see that the MCP flux from charged pion production exceeds that of Drell-Yan production all the way up to the kinematic threshold of the 120 GeV beam. Interestingly, this is true even for the conservative assumption of the target mass equal to the bare nucleon mass, while for dark photon production in Ref. Curtin et al. (2023), the relative benefit of charged pion scattering is somewhat smaller. As alluded to earlier, MadGraph becomes unreliable for small MCP masses as the virtuality of the photon becomes small (this can be seen in the spurious increase of the cross section for $m_{\chi}=400\ {\rm MeV}$ as $E_{\pi}\to 0$ in Fig. 4), so we do not show the flux for $m_{\chi}<400\ {\rm MeV}$. That said, based on the results of Ref. Curtin et al. (2023) for similar kinematics, we strongly expect that the flux will be subdominant to meson decays in this mass range and will not yield any additional gains in sensitivity.
Figure 5: Number of MCP $N_{\chi}$ incident at the DUNE ND, for $10^{21}$ POT and $\epsilon=10^{-3}$. Charged pion scattering exceeds Drell-Yan scattering for $m_{\chi}\gtrsim 1\ {\rm GeV}$ up to highest masses ($\sim 5\ {\rm GeV}$) allowed by the chiral cutoff, and is the dominant production mode between 2–3 GeV.